A total of 346 first year students of CMU are currently enrolled in the subject GEC15. If there are 26 more girls than boys. How many girls are there?


1
Expert's answer
2021-10-25T19:10:22-0400

Let x=x= the number of boys. Then there are x+26x+26 girls.


x+x+26=346x+x+26=346

2x=3202x=320

x=160x=160

160+26=186160+26=186

Then there are 186186 girls.